Suppose that Eleanor has a utility function given by: U=2x +4y. If the price of X is $10 then which of the following prices of Y is the highest price that will get her to still purchase units of Good X?

a. PY= $1

b. PY= $4

c. PY= $25

d. PY= $50

Please show work
Answer #1

Utility: U = 2x+4y

MRS = MUx/MUy

MRS = 2/4 = ½

Price ratio = Px/Py

The consumer will consume good x if MRS => Px/Py

½ =>10/Py

Py=>20

So, the correct answer is Py = $50.

The correct option is d) .